OpenStudy (anonymous):

Solve using the substitution method. 1/3x-1/5y= 7 1/6x+2/3= -4

OpenStudy (anonymous):

can u show how u got that plz

OpenStudy (anonymous):

Is the fraction (1/3)x or 1/(3x) ?

OpenStudy (anonymous):

my answer is this but im missing Y 1/3x-1/5y= 7 1/6x+2/3= -4 y = (-15/2)/ (23/30) y = -15(30)/2(23) = -15(15)/(23) = -225/23

OpenStudy (anonymous):

I just need the right answer

OpenStudy (anonymous):

i think its this 1/6x+2/3=-4 1/6x=-4-2/3 1/6x=-14/3 x=-28 substitute x=-28 into 1) 1/3(-28)-1/5y=7 -28/3+1/5y=7 1/5y=7+28/3 1/5y=49/3 y=245/3 or 81.666666666666666666666666666666...

OpenStudy (anonymous):

umm